Journalist Report – March 8th

Sol 24 – Panic at the Aquaponics

“A toast to all of you; it’s good to be with friends again.”

– Chapter 24 of The Martian Chronicles by Ray Bradbury

The shuttle landed in the night, not far from the Mars Desert Research Station. Alice, Corentin and I stepped out, treading lightly as the gentle Martian gravity welcomed us back. Light burst from the main airlock window; our crew was waiting for us! Waving hands and smiling faces were gathered behind the second airlock door, waiting for the pressure to equalize. When Jérémy let us in, I felt like I was home. I would have hugged them all, if I had had two functioning arms! The crew was reunited once and for all, ready to finish our mission together.

This morning, like every morning, at the beginning of the daily workout, the song Cheerleader came on, and though I desperately wanted to join in, I simply tapped my feet, and smiled at the thought that for the rest of the mission, I could be my team’s very own cheerleader!

Almost everything feels like it is going back to normal: after a few strange and chaotic Sols, we are all managing to ease back into our routine: taking our tests, handling data for our experiments, writing our reports. This makes me think that I have not yet mentioned what has been taking up most of Alice’s time these past few Sols. As crew Scientist, she is responsible for centralizing all the data our experiments produce, from images captured during EchoFinder sessions, to the many daily surveys each crewmember takes. To put it her way: “I don’t do data handling. I do big data handling!”
Throughout the day, we continued testing AI4U, prompting it with an emergency protocol written by Alexandre and Quentin which we aim to start simulating tomorrow: at a random time during the day, an alarm will go off, indicating the depressurization of one of the station’s modules. We will then test whether the AI is useful in getting us organized and helping us execute the emergency procedure!

Yesterday, Jérémy wrote that despite the circumstances, routine operations that keep the station up and running have never ceased taking place. Most notably, Adrien has continuously taken good care of the GreenHab. Unfortunately, for unknown reasons, a second fish has ascended to fish-heaven today. Out of the eight fishes originally brought to the MDRS by Adrien, we have lost both Mercury and Jupiter. Our botanist’s hypotheses range from a sudden increase in pH, to the possible emergence of bacteria, to the increasing temperatures in the GreenHab. The investigation is ongoing…

Although accumulated stress, our lack of sleep, and general state of fatigue are starting to take their toll on the crew after more than three weeks on Mars, surprisingly, one thing comes up more and more often in conversations: not wanting to leave…

GreenHab Report – March 8th

GreenHab Officer: Adrien Tison

Environmental control: Heater

Average Temperature: 74.4°F

Average humidity: 20%

Hours of supplemental light: 0 hour

Daily water usage for crops: 9.51 gallons

Daily water usage for research and/or other purposes: 2.11gallons

Water in the blue tank: 62.37 gallons

Time(s) of watering for crops: 9:30

Changes to crops: N/A

Narrative: I start feeling like this week really is testing us right now…

Today morning another fish was found dead in the fish tank. I immediately started to transfer them to a white ‘bucket’ with clear water (taking my time nonetheless because they needed to get used to this new water).

Then I conducted water tests to understand, and the values obtained were quite usual, apart from the pH which went a little bit high too quickly. It means that the kH is high, and it may be the reason.

From now on I will not put them back in the aquaponics system. Indeed, if it isn’t due to the pH, the water may be infected. Not being able to treat the water or change the entirety of it considering the amount of time left, I’d rather keep them safe and finish the mission and my experiment with a hydroponics system.

Apart from that, I watered and maintained plants and crops. According to what I see I think the next crews will be able to harvest and eat well.

I really love seeing how cucumber plants are thriving and growing more and more.

Harvest: 427 g of tomatoes, 5 g of sage, 2 g of thyme

Support/supplies needed: Donald, considering how the fish are doing right now, you can come to take the fish if you want. I do not know if it is possible, but I mean, Safety first for them. As I will not reintegrate them into the aquaponics system I thought it wouldn’t be a problem for you to take them if you want to.

Operations Report – March 8th

Crew 275 Operations Report 08Mar2023
SOL: 24

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet
• Spacesuit n°9
• Robotic observatory

Notes on non-nominal systems: None.

Spirit rover used: No
• Hours: 209.9
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Opportunity rover used: No
• Hours: 113.9
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Curiosity rover used: No
• Hours: 221.8
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Perseverance rover used: No
• Hours: 255.9
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

General notes on rovers: Nothing to report.

Summary of Hab operations: The Crew conducted routine operations.
• WATER USE: 13.39
• Water: 341.96 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: No

Summary of internet: Starlink (MDRSCampus-Guest) is used during COMMS and during the day for emails and to handle the medical emergency. MDRScampus is used for experimental purpose.

Summary of suits and radios: None.

Summary of GreenHab operations: The Crew botanist conducted routine operations. He transferred the fish to a new tank.
• WATER USE: 11.62 gallons
• Heater: On
• Supplemental light: No
• Harvest: 427 g of tomatoes – 5 g of sage – 2 g of thyme

Summary of ScienceDome operations: The Commander conducted experimental protocols.
• Dual split: Heat / From 10 pm to 7 am

Summary of RAM operations: No operations were conducted in the RAM today.

Summary of any observatory issues: The robotic observatory is non-nominal.

Summary of health and safety issues: None.

Questions, concerns, and requests to Mission Support:
• The lamp located on the right when entering the ScienceDome is missing a light bulb.

Sol Summary Report – March 8th

Crew 275 Sol Summary Report 08Mar2023

Sol: 24

Summary Title: Panic at the aquaponics

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Yesterday, we were very lucky that the surgery of the injured crewmember could take place so quickly. Everyone was back at the Hab before midnight, and we spent some time debriefing our respective days. Even if we are happy to all be back in simulation. We can all feel that we are exhausted by the 24 days of the mission and all the events that have happened in the meantime.

To make matters worse, another fish has been found floating dead in the aquarium today. The indices of water quality are still normal, except for a slightly high pH, so we do not really understand what happened. The remaining fish have been placed in another bucket, with new water and separated from the hydroponics system. We all hope that this will solve the problem. Since the fishes are no longer used to grow the plants and they are now in a relatively small bucket, maybe Donald would like to retrieve them before the end of the mission. The situation is far from ideal for them at the moment.

As the end of the mission is close, we have trained for one of our last experiments: the emergency protocols assisted by AI. In particular, we have rehearsed a protocol aimed at repairing a fictive hole causing a depressurization. Starting now, an alarm may ring anytime and any day between 14:00 and 18:00. When the alarm rings, we will have to proceed with the rehearsed protocol.

Look Ahead Plan:

We would like to perform an exploration EVA to the Special Region tomorrow (Sol 25), and another EVA dedicated to the disassembling of the atmospheric instruments on Friday (Sol 26).

Anomalies in work: None.

Weather: Cloudy, low -3°C/27°F, high 9°C/49°F.

Crew Physical Status: One crewmember recovering from surgery after a broken collarbone.

EVA: None.

Reports to be filed: Sol Summary, EVA Request, Operations Report, Journalist Report, GreenHab Report, HSO Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site